OpenStudy (anonymous):

I think the answer is - 1/5, but I need an A so I need to be positive. Heres the question: What is the slope of the line containing the ordered pair (2, 3) and (-4, 0)?

hartnn (hartnn):

its not -1/5 The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

m = - 4 - 3/ 0 - 2 = 7/ - 2 that's what I got now.

hartnn (hartnn):

no... x1 =2, y1 =3 x2=-4, y2=0

OpenStudy (anonymous):

m = 3 - 0/ 2 - 4 = 3/2?

hartnn (hartnn):

no...

hartnn (hartnn):

\(\huge m=\frac{y_1-y_2}{x_1-x_2}=\frac{3-0}{2-(-4)}=\frac{1}{2}\)

OpenStudy (anonymous):

Oh, that makes sense, now I feel stupid. Thanks for your help.

hartnn (hartnn):

welcome ^_^